The equation of a certain circle is x+y+6x-4y+8=0
i) find the equation of a tangent from the point (-1,5)
ii) the length of the tangent

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

The equation of a certain circle is x+y+6x-4y+8=0
i) find the equation of a tangent from the point (-1,5)
ii) the length of the tangent
-----------------
That's not a circle, it's a straight line.

Answer by richwmiller(17219)

x+y+6x-4y+8=0
is a straight line not a circle.
perhaps you mean
x^2+y^2+6x-4y+8=0
but (-1,5) is not on that circle